OpenStudy (anonymous):

Solve. 4 - x^2 over x^2 + 2x - 8.

OpenStudy (diyadiya):

Factor both of them 4 - x^2 and x^2 + 2x - 8.

OpenStudy (anonymous):

Okay? Then what?

OpenStudy (diyadiya):

You'll be able to cancel out common terms can you tell me what you got after factoring ? 4-x^2= ? x^2+ 2x - 8 = ?

OpenStudy (anonymous):

well, the common factor for 4 - x^2 is 2x, right?

Parth (parthkohli):

@fallingangel No, 4 - x^2 is prime.

OpenStudy (anonymous):

Oh... See? I need help. :/

OpenStudy (diyadiya):

No factor it .. 4-x^2 can be written as (2)^2 -(x)^2 \[a^2-b^2=(a+b)(a-b)\]

OpenStudy (diyadiya):

\[(2)^2-(x)^2=(2-x)(2+x)\]

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow x^2 -2x + 4x - 8 }\) \(\Large \color{purple}{\rightarrow x(x - 2) + 4(x - 2) }\) \(\Large \color{purple}{\rightarrow (4 + x)(x - 2) }\) ^^ FACTORIZATION

OpenStudy (diyadiya):

\[\Large \frac{4-x^2}{x^2+2x-8}=\frac{(4+x)(4-x)}{(4+x)(x-2)}\]

OpenStudy (diyadiya):

Now you can cancel out 4+x

OpenStudy (diyadiya):

@fallingangel did you understnad ?

OpenStudy (anonymous):

Thank you so much. :)

OpenStudy (anonymous):

Yes, I do.

Parth (parthkohli):

Now, just cross multiply the numerator of the first fraction and denominator of the other and vice-versa

OpenStudy (diyadiya):

@ParthKohli cross multiply ?

OpenStudy (anonymous):

That's what I was wondering.

Parth (parthkohli):

I mean, cross-multiply on both sides of the equation.

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow {a \over b} = {c \over d} \rightarrow ad = bc }\)

Parth (parthkohli):

That's what I meant.

OpenStudy (anonymous):

Where would I have used that though?

OpenStudy (diyadiya):

@ParthKohli you probably got the question wrong

Parth (parthkohli):

No wait, I thought it was an equation from one of your replies :P